THE HIGH RESOLUTION S_{1}←S_{0} SPECTRUM OF THE ANISOLE-WATER COMPLEX: WATER AS AN ACID. ^{a}

\underline{JASON W. RIBBLETT}, Department of Chemistry, Ball State University, Muncie IN 47302; DAVID R. BORST, and DAVID W. PRATT, Department of Chemistry, University of Pittsburgh, Pittsburgh PA 15260.

The rotationally resolved S_{1}←S_{0} electronic spectra of anisole and the 1:1 anisole-water complex have been recorded in a molecular beam. This is the first rotationally resolved electronic spectrum recorded in which water acts as an acid and the chromophore acts as base. The spectrum is split into two subbands separated by 730 MHz, having an intensity ratio of 3:1. This indicates that the water hydrogens are exchanged by a low barrier tunneling motion. The nature of this motion will be discussed.

aWork supported by NSF.